# Asymptotic behavior of the ground state energy of a fermionic Fröhlich multipolaron in the strong coupling limit

### v2

@@ 3 UV cut-off and partition into block modes
We will prove the inequality after adding an arbitrary $\epsilon>0$ on the right hand side. We consider a normalized $n_{j}$-fermion state $\psi_{j}$ with support in $B_{j}$ such that $E_{PT}^{(n_{j},\alpha)}(\psi_{j})<E_{PT}^{(n_{j},\alpha)}(B_{j})+\frac{\epsilon}{m}$. Let $\psi_{1}\wedge\ldots\wedge\psi_{m}$ be the normalized anti-symmetrized tensor product of $\psi_{1},\dots,\psi_{m}$. Then, since the supports of $\psi_{1},\dots,\psi_{m}$ are disjoint, we obtain $\rho_{\psi_{1}\wedge\ldots\wedge\psi_{m}}=\sum_{j=1}^{m}\rho_{\psi_{j}}$. Moreover, for every local operator $A$ that is symmetric with respect to the coordinates $x_{1},\dots,x_{N}$ we have that
$\langle\psi_{1}\wedge\ldots\wedge\psi_{m},A\psi_{1}\wedge\ldots\wedge\psi_{m}\rangle=\langle\psi_{1}\otimes\ldots\otimes\psi_{m},A\psi_{1}\otimes\ldots\otimes\psi_{m}\rangle.$ (3.14)
As a consequence, decomposing the interaction terms to the ones that are internal within a cluster and interaction terms between different clusters we obtain
$E_{PT}^{(N,\alpha)}(\psi_{1}\wedge\ldots\wedge\psi_{m})\,$ $=\,\sum_{j=1}^{m}E_{PT}^{(n_{j},\alpha)}(\psi_{j})-\alpha\sum_{i\neq j}\int_{B_{i}}\int_{B_{j}}\frac{\rho_{\psi_{i}}(x)\rho_{\psi_{j}}(y)}{|x-y|}\,\mathrm{d}x\,\mathrm{d}y$ (3.15)
where $C_{i},C_{j}$ denote the $i$-th and $j$-th clusters, and the factor $\frac{1}{2}$ stems from replacing $\sum_{i<j}$ by $\sum_{i\neq j}$. Using the fact that $\rho_{\psi_{i}}(x)=|C_{i}|\int_{\mathbb{R}^{3(|C_{i}|-1)}}|\psi_{i}(x,X)|^{2}\,\mathrm{d}X$, we arrive at
$E_{PT}^{(N,\alpha)}\,$ $\leq\,\sum_{j=1}^{m}E_{PT}^{(n_{j},\alpha)}(\psi_{j})+(\nu-2)\alpha\sum_{i<j}\int_{B_{i}}\int_{B_{j}}\frac{\rho_{\psi_{i}}(x)\rho_{\psi_{j}}(y)}{|x-y|}\,\mathrm{d}x\,\mathrm{d}y\,.$ (3.17)
Since $|x-y|\geq R$ for $x\in B_{i}$, $y\in B_{j}$, $i\neq j$, we obtain that
$E_{PT}^{(N,\alpha)}\,\leq\,\sum_{j=1}^{m}E_{PT}^{(n_{j},\alpha)}(\psi_{j})+\frac{(\nu-2)_{+}\alpha N^{2}}{2R}\,$, (3.18)
Using the last equality together with $E_{PT}^{(n_{j},\alpha)}(\psi_{j})<E_{PT}^{(n_{j},\alpha)}(B_{j})+\frac{\epsilon}{m}$, we arrive at the desired result. ∎
@@ 4 Proof of main result
We are finally ready to prove Theorem 1.1.
@@
Employing Lemma 2.2, Proposition 2.5 and Remark 2.6, implies that
@@
with the notation of Lemma 2.2. Recalling $\beta_{j}=1-\frac{8n_{j}\alpha}{\pi\Lambda_{j}}$, and collecting Lemma 3.1 and Propositions 3.2, and 3.3, we obtain
@@
Employing Lemma 3.6 and using the facts that $1-\beta_{j}=\frac{8n_{j}\alpha}{\pi\Lambda_{j}}$ and $R_{j}=\frac{1}{2}(3n_{j}-1)R\leq\frac{3}{2}n_{j}R$, we obtain that
@@
for some constant $C$ dependent on $A$, $V$, and $|\nu|$ only. This completes the proof of our main Theorem 1.1. ∎
@@ Appendix A Derivation of the Pekar-Tomasevich functional
$E_{PT}^{(N,\alpha)}(\Psi_{N})\,$ $=\,\sum_{j=1}^{N}\left\langle\Psi_{N},\Big(D_{A_{\alpha},x_{j}}^{2}+V_{\alpha}(x_{j})+\sum_{k=j+1}^{N}\frac{\nu\alpha}{|x_{j}-x_{k}|}\Big)\Psi_{N}\right\rangle$ (A.1)
@@
We have for $\|\eta\|_{\mathcal{F}}=\|\Psi_{N}\|_{\bigwedge_{j=1}^{N}L^{2}(\mathbb{R}^{3})}=1$
@@
Appendix B Localization results
@@
As in [Wel15], we start by subdividing $\mathbb{R}^{3}$ into the ’area of influence’ of the single multipolarons. More precisely, we define
@@
The first term in (B.5) corresponds to the localized phonon field energy, and the second to the localized electron-phonon coupling. The last term in (B.5) corresponds to the interactions with all other clusters. To absorb it, we conjugate with the unitary Weyl transform
@@
The following result is taken directly from [Wel15, Lemma 3.4].
B.1 Lemma.
Let $\Phi_{0}$ be as constructed in Lemma 2.2, and recall $F_{1}$ and $F_{2}$ from (B.12). Then we have
  1. (i)
@@
For the proof we refer to [Wel15, Lemma 3.4]. The proof of part (ii) relies on the known identity
@@ Appendix C Phonon integration
We now provide the proofs of [LT97] with their proper adaption to the multipolaron setting, see also [Wel15]. The reason for which we present these proofs is to guarantee that the localized electronic support inside of a given ball can, in fact, be preserved.
@@
We start by shifting the center $x_{0}$ of $B$ to the origin. For that, we introduce the total momentum
@@
and the center of $B$ by the origin. Since the ground state energy is invariant under conjugation with unitary transformations, it suffices to show the statement for the shifted external potentials.
@@
where both, $H^{(n,\alpha)}_{\Lambda}$ and $H^{(n,\alpha,\beta)}_{\mathrm{block}}$, contain the shifted fields $V_{\alpha}$, $A_{\alpha}$. Since $k_{\ell}$ was chosen arbitrarily, employing (C.9) and (C.6) now implies
@@
C.1 Lemma.
Recall that $\beta=1-\frac{4n\alpha}{\pi\Lambda}$. There exists a constant $C>0$ dependent on $A$, $V$, and $|\nu|$ s.t.
@@
Building upon ideas in [Wel15], we reduce $E_{PT}^{(n,\alpha,\beta^{-1})}(B)$ to $E_{PT}^{(n,\alpha)}(B)$ by exploiting concavity of $\lambda\mapsto E_{PT}^{(n,\alpha,\lambda)}(B)$. More precisely, as an infimum over concave functions, $\lambda\mapsto E_{PT}^{(n,\alpha,\lambda)}(B)$ itself is concave. In particular, we obtain that for $\beta=1-\frac{4n\alpha}{\pi\Lambda}$ and $\frac{\Lambda}{n\alpha}$ large enough s.t. $\beta^{-1}\in(1,2)$
@@
To complete the proof, it suffices to establish upper bound on $E_{PT}^{(n,\alpha,\lambda)}$. A scaling argument analogous to Remark 1.3 yields that
@@
Hardy’s inequality and the diamagnetic inequality [LL96, Theorem 7.21], together with the infinitesimal form-boundedness of $V$ w.r.t. $\Delta$, and $A_{k}\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{3})$ imply the general bound
@@
for some constant $C>0$ dependent on AA,$V$ and $|\nu|$. In the case $\nu>2$ and in the absence of external fields $A$, $V$, [FLST11, Theorem 3] implies
@@
for some $C=C(\nu)>0$, where $\lim_{\nu\to 2^{+}}C(\nu)\to\infty$. Their result holds independently of statistics of particles. The case $\nu=2$ without external fields is covered in [GMl10] and establishes the stability bound (C.21) in the fermionic case. Together with (C.18), we conclude the proof. ∎
@@
$h_{z}$ $=\sum_{j=1}^{n}\Big(D_{A_{\alpha},x_{j}}^{2}+\beta^{-1}V_{\alpha}(x_{j})+\frac{\sqrt{\alpha}}{\sqrt{2}\pi\beta}\sum_{\ell\in\Lambda_{P}}M_{\ell}\left(z_{\ell}e^{ik_{\ell}x_{j}}+\overline{z}_{\ell}e^{-ik_{\ell}x_{j}}\right)+\sum_{k=j+1}^{N}\frac{\nu\alpha\beta^{-1}}{|x_{j}-x_{k}|}\Big)$
$\qquad+\sum_{\ell\in\Lambda_{P}}(|z_{\ell}|^{2}-1).$

### v3

@@ 3.2 Cluster synthesis
Consider a normalized $n_{j}$-fermion state $\psi_{j}$ with support in $B_{j}$ such that $\mathcal{E}_{PT}^{(n_{j},\alpha)}(\psi_{j})<E_{PT}^{(n_{j},\alpha)}(B_{j})+\frac{\epsilon}{m}$. Let $\psi_{1}\wedge\ldots\wedge\psi_{m}$ be the normalized anti-symmetrized tensor product of $\psi_{1},\dots,\psi_{m}$. Since the supports of $\rho_{\psi_{1}},\dots,\rho_{\psi_{m}}$ are disjoint, Lemma 3.5 implies that
$E_{PT}^{(N,\alpha)}\,\leq\,\sum_{j=1}^{m}\mathcal{E}_{PT}^{(n_{j},\alpha)}(\psi_{j})+\frac{(\nu-2)_{+}\alpha N^{2}}{2R}\,$, (3.21)
Using the last equality together with $\mathcal{E}_{PT}^{(n_{j},\alpha)}(\psi_{j})<E_{PT}^{(n_{j},\alpha)}(B_{j})+\frac{\epsilon}{m}$, and sending $\varepsilon\to 0$, we arrive at the desired result. ∎
@@ 4 Proof of main result
With the localization and strong-coupling estimates established, we are now in position to prove Theorem 1.2.
@@
Employing Prop. 2.2, Proposition 2.5 and Remark 2.6, implies that
@@
with the notation of Prop. 2.2. Recalling $\beta_{j}=1-\frac{8n_{j}\alpha}{\pi\Lambda_{j}}$, and collecting Lemma 3.1 and Propositions 3.2, and 3.3, we obtain
@@
We will choose $\Lambda_{j}$ s.t. $\alpha/\Lambda_{j}\to 0$ as $\alpha\to\infty$. Then let $\alpha$ be sufficiently large such that $1-\beta_{j}=\frac{8n_{j}\alpha}{\pi\Lambda_{j}}\in(0,\frac{1}{2})$. Employing Corollary 3.7 and using the fact that for $R_{j}=\frac{1}{2}(3n_{j}-1)R\leq\frac{3}{2}n_{j}R$, we obtain that
@@
for some constant $C$ dependent on $A$, $V$, and $|\nu|$, as clarified in Remark 1.3; this dependence arises from the bound established in Lemma C.1 in the appendix. This completes the proof of our main Theorem 1.2. ∎
@@ A.1 Pekar’s argument
Notice that for $\|\eta\|_{\mathcal{F}}=\|\Psi_{N}\|_{\bigwedge_{j=1}^{N}L^{2}(\mathbb{R}^{3})}=1$
@@ A.2 Sketch of rigorous version of Pekar’s argument
Appendix B Localization arguments
Proof of Lemma 2.1.
Assume the statement holds true for $N$. Consider an additional point $y_{N+1}\in\mathbb{R}^{3}$. If $\operatorname{dist}(B_{R}(y_{N+1}),B_{j})\geq R$ for all $j\in\{1,\ldots,m\}$, define $B_{m+1}:=B_{R}(y_{N+1})$ and $C_{m+1}:=\{N+1\}$.
@@ Appendix B Localization arguments
As in [42], we start by subdividing $\mathbb{R}^{3}$ into the ’area of influence’ of the single multipolarons. More precisely, we define
@@
The first term in (B.5) corresponds to the localized phonon field energy, and the second to the localized electron-phonon coupling. The last term in (B.5) corresponds to the interactions with all other clusters. To absorb the latter, we conjugate the Hamiltonian with the unitary Weyl transform
@@
The following result is taken directly from [42, Lemma 3.4].
B.1 Lemma (Inter-cluster interaction bounds).
Let $\Phi_{0}$ be as constructed in Prop. 2.2, and recall $F_{1}$ and $F_{2}$ from (B.12). Then we have
  1. 1.
@@
For the proof we refer to [42, Lemma 3.4]. The proof of part (ii) relies on the known identity
@@ Appendix C Phonon integration
We now provide the proofs of [29] with their proper adaption to the multipolaron setting, see also [42]. The reason, for which we present these proofs, is to guarantee that the localized electronic support inside of a given ball can, in fact, be preserved.
@@
We start by shifting the center $x_{0}$ of $B$ to the origin. To that end, we introduce the total momentum
@@
and the center of $B$ by the origin. Since the ground-state energy is invariant under conjugation with unitary transformations, it suffices to show the statement for the shifted external potentials.
@@
where both, $H^{(n,\alpha)}_{\Lambda}$ and $H^{(n,\alpha,\beta)}_{\mathrm{block}}$, contain the shifted fields $V_{\alpha}$, $A_{\alpha}$. Since $k_{\ell}$ was chosen arbitrarily, we employ (C.6) and (C.9) to bound the error terms. Modulo translating $H^{(n,\alpha)}_{\Lambda}$ and $H^{(n,\alpha,\beta)}_{\mathrm{block}}$ back, we thus obtain that
@@
C.1 Lemma (Concavity bound).
Assume that $\beta=1-\frac{4n\alpha}{\pi\Lambda}\in(\frac{1}{2},1)$. There exists a constant $C>0$ dependent on $A$, $V$, and $|\nu|$ s.t.
@@
Building upon ideas in [42], we reduce $E_{PT}^{(n,\alpha,\beta^{-1})}(B)$ to $E_{PT}^{(n,\alpha)}(B)$ by exploiting concavity of $\lambda\mapsto E_{PT}^{(n,\alpha,\lambda)}(B)$. More precisely, as an infimum over concave functions, $\lambda\mapsto E_{PT}^{(n,\alpha,\lambda)}(B)$ itself is concave. In particular, since $\beta^{-1}\in(1,2)$, we obtain that
@@
To complete the proof, it suffices to establish an upper bound on $E_{PT}^{(n,\alpha,\lambda)}$. A scaling argument analogous to Remark 1.1 yields that
@@
Hardy’s inequality and the diamagnetic inequality [27, Theorem 7.21], together with the infinitesimal form-boundedness of $V$ w.r.t. $\Delta$, and $A_{k}\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{3})$ imply the general bound
@@
for some constant $C>0$ dependent on AA,$V$ and $|\nu|$. In the case $\nu>2$ and in the absence of external fields $A$, $V$, [15, Theorem 3] implies
@@
for some $C=C(\nu)>0$, where $\lim_{\nu\to 2^{+}}C(\nu)\to\infty$. Their result holds independently of statistics of particles. The case $\nu=2$ without external fields is covered in [18] and establishes the stability bound (C.21) in the fermionic case. Together with (C.18), we conclude the proof. ∎
@@
$h_{z}$ $=\sum_{j=1}^{n}\Big(D_{A_{\alpha},x_{j}}^{2}+\beta^{-1}V_{\alpha}(x_{j})+\frac{\sqrt{\alpha}}{\sqrt{2}\pi\beta}\sum_{\ell\in\Lambda_{P}}M_{\ell}\left(z_{\ell}e^{ik_{\ell}x_{j}}+\overline{z}_{\ell}e^{-ik_{\ell}x_{j}}\right)$
$\qquad+\sum_{k=j+1}^{N}\frac{\nu\alpha\beta^{-1}}{|x_{j}-x_{k}|}\Big)+\sum_{\ell\in\Lambda_{P}}(|z_{\ell}|^{2}-1).$